A 2.32 m long pole is balanced vertically on its tip. It starts to fall and its lower end does not slip. What will be the speed of the upper end of the pole just before it hits the ground?
Answer #1

mgL/2 = 0.5 I ω2

ω2 = mgL/I

ω2 = mgL/(mL2/3)

     = 3*g/L

     = 3*9.8/2.32

     = 12.67241

ω = 3.5598

v = L ω = 2.32 * 3.5598 = 8.259 m/s